# Directionality and quantum backfire in continuous-time quantum walks from delocalized states: Exact results

### v1

@@ Abstract
We derive analytical results for continuous-time quantum walks from a new class of initial states with tunable delocalization. The dynamics are governed by a Hamiltonian with complex hopping amplitudes. We provide closed-form equations for key observables, revealing three notable findings: (1) the emergence of directed quantum transport from completely unbiased initial conditions; (2) a quantum backfire effect, where greater initial delocalization enhances short-time spreading but counterintuitively induces a comparatively smaller long-time spreading after a crossing time $t_{\mathrm{cross}}$; and (3) an exact characterization of survival probability, showing that the transition to an enhanced $t^{-3}$ decay is a fine-tuned effect. Our work establishes a comprehensive framework for controlling quantum transport through the interplay between intermediate initial delocalization and Hamiltonian phase.
@@
I Introduction
Quantum walks (QWs) are powerful frameworks for modeling quantum transport and developing quantum algorithms [1, 2, 3, 4, 5]. Their framework encompasses both discrete-time (DTQWs [6]) and continuous-time (CTQWs [7]) variants. Conventional versions of the DTQW and CTQW exhibit remarkable properties such as ballistic spread and bimodal distributions, but a richer phenomenology emerges with nonstandard versions of these models [8, 9, 10, 11, 12, 13]. QWs can be simulated from local states or extended (also called nonlocal or delocalized) states  [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
The first studies with an explicit focus on QWs from nonlocal initial conditions emerged in 2006, with key contributions from Refs. [14, 15, 16, 17]. In [14] the authors showed that delocalized initial conditions tend to enhance the entanglement between the internal (spin) and external (position) degrees of freedom in DTQWs. The authors of [15] analyzed the scaling of the survival probability of DTQWs with initial delocalized states. In turn, [16, 17] studied CTQW and DTQWs from extended initial states. The ability to tailor the propagation of QWs through extended initial conditions was systematically shown in [18]. The use of QWs from a Gaussian initial state to generate cat states was reported in [21]. In [29], DTQWs with a fully delocalized initial state were studied in the context of rogue waves. It was shown in [30] that both CTQWs and DTQWs from a delocalized state can serve as platforms for state transfer. The first explicit experimental demonstration of QWs from initial superposition states was presented in [31].
Recently, Ref. [32] investigated CTQWs initiated from both fully localized ($|0\rangle$) and fully delocalized ($|\pm k\rangle$) states. While these extreme cases have been thoroughly examined in the literature, the intermediate regime remains largely unexplored. Our work addresses this gap.
II Model
@@
Unlike all the works mentioned in the previous section, we define a new class of tunable delocalized initial states:
@@
III Results
III.1 Wavefunction
Figure 1: Probability distributions for $\alpha=\pi/2$ at $\gamma t=50$. Our results show that while both localized ($D=0$, bottom) and fully delocalized ($D=1$, top) initial states yield symmetric spreading, intermediate delocalization ($D=0.5$, middle) generates a pronounced bias. The analytical results were obtained with Eq. (7) and $P(x,t)=\left|\psi(x,t)\right|^{2}$.
Figure 2: Absolute value of the average group velocity $|\absolutevalue{\langle v_{g}\rangle/\gamma}$, Eq. (9), as a function of the delocalization parameter $D$ for various phases $\alpha$. While the extreme states ($D=0$ and $D=1$) yield zero net velocity for any $\alpha$, a maximum bias emerges at intermediate delocalization ($D=0.5$), demonstrating tunable directed transport.
Due to translational symmetry, the Hamiltonian is diagonal in the momentum basis. The momentum eigenstates $|k\rangle$ are Fourier transforms of the position basis states $|x\rangle$:
@@ III.1 Wavefunction
where the quasimomentum $k$ is in the first Brillouin zone, $k\in[-\pi,\pi]$. Applying the Hamiltonian to $|k\rangle$ yields the energy dispersion relation: $E(k)=-2\gamma\cos(\alpha-k).$
@@
Using the Jacobi-Anger expansion to evaluate this integral, the final form of the wavefunction is found to be:
$\psi(x,t)=\sqrt{1-D}\tilde{J}_{x}(2\gamma t)\\ +\sqrt{\frac{D}{2}}\left(e^{\mathrm{i}\alpha}\tilde{J}_{x-1}(2\gamma t)+e^{-\mathrm{i}\alpha}\tilde{J}_{x+1}(2\gamma t)\right),$ (7)
@@
Figure 1 shows the analytically obtained probability distributions, which are in excellent agreement with the numerical results. For the phase value considered ($\alpha=0.5$), the localized ($D=0$) and fully delocalized ($D=1$) initial setups, which are commonly considered in the literature, both exhibit symmetric distributions during the system’s evolution. The fully delocalized case leads to a lower probability in the central region. In contrast, the intermediate delocalization ($D=0.5$) leads to a biased spreading, despite the same phase factor value and an initially symmetric state. To further analyze this asymmetrical spreading, we calculate the average position as a function of these parameters.
III.2 Average Position
The average position $\langle X\rangle(t)$ is given by Ehrenfest’s theorem, $d\langle X\rangle/dt=\langle v_{g}\rangle(t)$, where $v_{g}(k)=\partial E(k)/\partial k=2\gamma\sin(\alpha-k)$ is the group velocity. The expectation value of the group velocity is
$\langle v_{g}\rangle(t)=\int_{-\pi}^{\pi}|\psi(k,t)|^{2}v_{g}(k)dk.$ (8)
Since the momentum distribution is time independent, $|\psi(k,t)|^{2}=|\psi(k,0)|^{2}$, the average group velocity $\langle v_{g}\rangle$ is constant.
The initial state distribution $|\psi(k,0)|^{2}$ is an even function of $k$. The group velocity $v_{g}(k)=2\gamma(\sin k\cos\alpha-\cos k\sin\alpha)$ has odd and even parts. Only its even part, $2\gamma\sin\alpha\cos k$, contributes to the integral, leading to
@@ III.2 Average Position
A net drift, characterized by $\langle v_{g}\rangle\neq 0$, emerges from the interplay between the initial state’s delocalization ($D$) and the Hamiltonian phase ($\alpha$). The drift vanishes for the extreme cases $D=0$ and $D=1$ but is maximized for intermediate delocalization, as shown in Fig. 2. Given that the initial average position is zero, $\langle X\rangle(0)=0$, The average position is then simply $\langle X\rangle(t)=\langle v_{g}\rangle t.$
Figure 3: Time evolution of the mean square displacement (MSD) obtained with Eq. (12). The top panel ($\alpha=0$) shows the quantum backfire effect: (i) for $t<t_{\mathrm{cross}}$, a larger initial MSD(0) promotes a larger MSD($t$); (ii) for $t>t_{\mathrm{cross}}$, this relationship inverts: a larger initial MSD(0) produces a smaller MSD($t$). The dashed vertical line marks $\gamma t_{\mathrm{cross}}$. The bottom panel ($\alpha=\pi/2$) exhibits no-crossing behavior: The ordering of MSD curves with respect to $D$ is preserved for all time.
Figure 4: Dependence of the MSD crossing time $\gamma t_{\mathrm{cross}}$ on the phase $\alpha$, Eq. (13). Two regimes are shown. 1) No-crossing (yellow, $\sin^{2}\alpha\geq 1/2$): MSD curves remain ordered for all time. 2) Crossing (blue, $\sin^{2}\alpha<1/2$): MSD curves intersect at $t=t_{\mathrm{cross}}$. For $t>t_{\mathrm{cross}}$, the ordering inverts, demonstrating the quantum backfire effect where a greater initial delocalization ($D$) boosts short-time spreading, but is detrimental to the long-time propagation.
@@
III.3 Mean Square Displacement
The mean square displacement (MSD), defined as $\mathrm{MSD}(t)=\langle(X-X_{0})^{2}\rangle(t)$, measures the wavepacket’s spreading around its initial average position $X_{0}\equiv\langle X\rangle(0)$. For our initial state, symmetry dictates $\langle X\rangle(0)=0$, which simplifies the expression:
$\mathrm{MSD}(t)=\langle X^{2}\rangle(t)-\langle X\rangle(0)^{2}=\langle X^{2}\rangle(t).$ (10)
Using $X=\mathrm{i}\,\partial_{k}$ and $\psi(k,t)=e^{-\mathrm{i}E(k)t}\,\psi(k,0)$, we find that $\left|\mathrm{i}\,\partial_{k}\psi(k,t)\right|^{2}=(t\,\partial_{k}E)^{2}\,|\psi(k,0)|^{2}+\left|\partial_{k}\psi(k,0)\right|^{2}$, leading directly to $\langle X^{2}\rangle(t)=\langle X^{2}\rangle(0)+t^{2}\langle v_{g}^{2}\rangle$.
The initial MSD, $\langle X^{2}\rangle(0)=D$, quantifies the initial delocalization. This result is intuitive: it equals 0 for a state perfectly localized at the origin ($D=0$) and 1 for a state delocalized over sites $x=\pm 1$ ($D=1$).
The average squared group velocity is given by $\langle v_{g}^{2}\rangle=\int v_{g}(k)^{2}|\psi(k,0)|^{2}dk$. Substituting $v_{g}(k)=-2\gamma\sin(\alpha-k)$ and the expression for $|\psi(k,0)|^{2}$ from Eq. (4), we evaluate the integral to obtain:
$\langle v_{g}^{2}\rangle=2\gamma^{2}\left(1-\frac{D}{2}+D\sin^{2}\alpha\right).$ (11)
Combining these results yields the exact analytical expression for the MSD:
$\mathrm{MSD}(t)=D+2\gamma^{2}t^{2}\left(1-\frac{D}{2}+D\sin^{2}\alpha\right).$ (12)
@@ III.3 Mean Square Displacement
$\gamma t_{\mathrm{cross}}(\alpha)=\frac{1}{\sqrt{1-2\sin^{2}\alpha}}.$ (13)
@@
The existence of a finite crossing time reveals a regime of counterintuitive dynamics. For a fixed time $t>t_{\mathrm{cross}}(\alpha)$ within the interval $0<\alpha<\pi/4$, Eq. (12) shows that the MSD becomes a decreasing function of the delocalization parameter $D$. This means that for these times, a more delocalized initial state (higher $D$) results in a smaller spatial spread of the wavepacket, the opposite of the intuitively expected outcome. This counterintuitive phenomenon, where the strategy of increasing initial delocalization to enhance spreading instead weakens it after a characteristic time, is formally analogous to the “Backfire Effect” known in cognitive science [38]. There, corrective evidence can paradoxically strengthen a person’s misconceptions. Here, a more delocalized initial state improves the short-time spreading, but leads to worse asymptotic spreading performance for a proper combination of parameters. We therefore designate this as a Quantum Backfire Effect.
This effect is clearly visible in the post-crossing regime (blue shading in Fig. 4 for $\sin^{2}\alpha<1/2$), where the MSD for $D=1$ becomes lower than for $D=0.5$, which in turn becomes lower than for $D=0$, as shown in the top panel of Fig. 3. For times $t<t_{\mathrm{cross}}(\alpha)$, the intuitive ordering, where greater initial delocalization leads to greater MSD, still holds (see Fig. 3).
Figure 5: Time evolution of $P_{\text{surv}}$ on a log-log scale for $\alpha=\pi/2$, Eq. (15). The $D=1$ case exhibits enhanced decay ($\sim t^{-3}$), while any partially delocalized state ($D<1$) shows the standard scaling ($\sim t^{-1}$). Dashed lines show the analytical asymptotic predictions.
III.4 Survival Probability
@@
Starting with an initial condition where $P_{\text{surv}}(0)=1$, the wavepacket spreading leads to a decrease in this quantity. The decay analysis of this value informs how fast the particle leaves the region. Using the probability distribution we get:
@@
This result allows for a precise analysis of the asymptotic decay. For the generic case ($D\neq 1$), the long-time behavior is governed by the squared Bessel function terms, $J_{n}(z)^{2}\approx 1/(\pi z)$, while the oscillatory cross-term $J_{0}J_{2}$ averages to zero. This leads to the decay $P_{\text{surv}}(t)\approx(3-2D\sin^{2}\alpha+D)/(2\pi\gamma t)$.
However, a exception occurs for the fine-tuned parameters $D=1$ and $\alpha=\pi/2+m\pi$ (where $m\in\mathbb{Z}$). In this case, Eq. (15) simplifies perfectly to $P_{\text{surv}}(t)=(J_{0}(2\gamma t)+J_{2}(2\gamma t))^{2}$. Using the Bessel function recurrence relation $J_{n-1}(z)+J_{n+1}(z)=(2n/z)J_{n}(z)$, this becomes: $P_{\text{surv}}(t)=J_{1}(2\gamma t)^{2}/(\gamma^{2}t^{2})$. For long times, this expression decays as: $P_{\text{surv}}(t)\approx 1/(\pi\gamma^{3}t^{3})$.
These analytical results show that the enhanced temporal decay $t^{-3}$ occurs only in the case $D=1$ for specific values of $\alpha$. Other combinations of $D$ and $\alpha$ lead to the usual scaling with $t^{-1}$. These findings are numerically confirmed in Fig. 5. Thus, we clarify that in a previous work [32] the breakdown of the standard scaling for $P_{\mathrm{surv}}(t)$ was a result of the special choice of the initial condition.
IV Final Remarks
We presented an analytical treatment for CTQWs with a tunable delocalization parameter $D$. Our main results are given by Eqs. (9,12,15). All these equations are associated with properties that correspond to experimentally accessible observables, making our proposal potentially testable in quantum platforms [45]. The key insights from our work are threefold.
First, we demonstrated an emerging directional spreading from unbiased initial states with intermediate delocalization, $0<D<1$. This result comes from the synergy between the initial state’s delocalization and a phase parameter intrinsic to the Hamiltonian.
Second, our analysis revealed a Quantum Backfire Effect: a phenomenon where an increase in the initial state’s delocalization can increase the short-time spreading, but can decrease the long-term spatial spread. This effect occurs after a crossing time, $t_{\mathrm{cross}}$.
Third, by computing the exact survival probability, we elucidated that the transition from a standard scaling $\sim t^{-1}$ to an enhanced decay $\sim t^{-3}$ is a fine-tuned effect that occurs for the specific case of a fully extended initial state ($D=1$) and for specific Hamiltonian phases.
The control of transport dynamics in QWs is an important topic in quantum information science [3, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Taking a broad perspective, our findings offer novel insights into how the interplay between complex hopping and delocalized initial states can lead to surprising phenomena related to transport properties.

### v2

@@ Abstract
We derive analytical results for continuous-time quantum walks from a class of initial states with tunable delocalization. The dynamics are governed by a Hamiltonian with complex hopping amplitudes. We provide closed-form equations for key observables, revealing three notable findings: (1) the emergence of directed quantum transport from completely unbiased initial conditions; (2) a quantum backfire effect, where greater initial delocalization enhances short-time spreading but counterintuitively induces a comparatively smaller long-time spreading after a crossing time $t_{\mathrm{cross}}$; and (3) an exact characterization of survival probability, showing that the transition to an enhanced $t^{-3}$ decay is a fine-tuned effect. Our work establishes a comprehensive framework for controlling quantum transport through the interplay between intermediate initial delocalization and Hamiltonian phase.
@@
Introduction.
Quantum walks (QWs) serve as versatile platforms for a wide range of fields, spanning from quantum algorithms [47, 26, 36, 28, 16] to the simulation of physical phenomena such as topological phases [62], relativistic effects [55], two-dimensional transport phenomena [57], and laser physics [22]. Their applicability extends even to interdisciplinary areas such as finance [14] and complex networks [9].
The framework of QWs encompasses both discrete-time (DTQWs [5]) and continuous-time (CTQWs [18]) variants. Conventional versions of these models exhibit remarkable properties such as ballistic spread and bimodal distributions, but a richer phenomenology emerges in nonstandard scenarios [45, 44, 53, 49, 67, 63]. A key degree of freedom in these systems is the initial condition: QWs can be initiated from local states or extended (also called nonlocal or delocalized) states [4, 3, 55, 54, 15, 6, 32, 66, 41, 43, 42, 19, 33, 58, 38, 10, 17, 56, 12].
First studies on nonlocal initial conditions appeared in 2006 [4, 3, 55, 54], revealing that delocalization can enhance entanglement between internal (spin) and external (position) degrees of freedom [4] and modify the scaling of survival probabilities [3]. Subsequent research demonstrated that extended initial states allow for the tailoring of wavepacket propagation [15], the generation of non-classical cat states [66], and the study of rogue waves [10]. Furthermore, delocalized states have been identified as effective resources for state transfer protocols [17], and their experimental feasibility has been successfully demonstrated using initial superposition states [56].
Recently, Ref. [12] investigated CTQWs initiated from both fully localized ($|0\rangle$) and fully delocalized ($|\pm x_{0}\rangle$) states. While these extreme cases have been thoroughly examined, the intermediate regime remains largely unexplored. Our work addresses this gap.
Model.
@@ Introduction.
Unlike all the works mentioned in the previous section, we define another class of tunable delocalized initial states:
@@
Figure 1: Probability distributions for $\alpha=\pi/2$ at $t=50$. Our results show that while both localized ($D=0$, bottom) and fully delocalized ($D=1$, top) initial states yield symmetric spreading, intermediate delocalization ($D=0.5$, middle) generates a pronounced bias. The analytical results were obtained using Eq. (7) and $P(x,t)=\left|\psi(x,t)\right|^{2}$.
Figure 2: Absolute value of the average group velocity $|\absolutevalue{\langle v_{g}\rangle}$ [Eq. (9)] as a function of the delocalization parameter $D$ for various phases $\alpha$. Solid lines represent the exact analytical results from Eq. (9), while symbols show numerical simulations. While the extreme states ($D=0$ and $D=1$) yield zero net velocity for any $\alpha$, a maximum bias emerges at intermediate delocalization ($D=0.5$), demonstrating tunable directed transport.
Results
Due to translational symmetry, the Hamiltonian is diagonal in the momentum basis. The momentum eigenstates $|k\rangle$ are Fourier transforms of the position basis states $|x\rangle$,
@@ Results
where the quasimomentum $k$ is in the first Brillouin zone, $k\in[-\pi,\pi]$. Applying the Hamiltonian to $|k\rangle$ yields the energy dispersion relation $E(k)=-2\gamma\cos(\alpha-k).$
@@
Using the Jacobi-Anger expansion to evaluate this integral (Sec. I of the Supplemental Material [2]), the final form of the wavefunction is found to be:
$\psi(x,t)=e^{\mathrm{i}x\alpha}\bigg[\sqrt{1-D}\,\tilde{J}_{x}(2\gamma t)\\ +\sqrt{\frac{D}{2}}\left(e^{\mathrm{i}\alpha}\tilde{J}_{x+1}(2\gamma t)+e^{-\mathrm{i}\alpha}\tilde{J}_{x-1}(2\gamma t)\right)\bigg],$ (7)
@@
Figure 1 shows the analytically obtained probability distributions, which are in excellent agreement with the numerical results. For the phase value considered ($\alpha=\pi/2$), the localized ($D=0$) and fully delocalized ($D=1$) initial setups, which are commonly considered in the literature, both exhibit symmetric distributions during the system’s evolution. The fully delocalized case leads to a lower probability in the central region. In contrast, the intermediate delocalization ($D=0.5$) leads to a biased spreading, despite the same phase factor value and an initially symmetric state. To further analyze this asymmetrical spreading, we calculate the average position as a function of these parameters.
We evaluate $\langle X\rangle(t)$ directly in momentum space using $X=\mathrm{i}\partial_{k}$ and $\psi(k,t)=e^{-\mathrm{i}E(k)t}\psi(k,0)$, which gives
$\langle X\rangle(t)=\int_{-\pi}^{\pi}\psi^{*}(k,0)\left[t\,v_{g}(k)+\mathrm{i}\partial_{k}\right]\psi(k,0)dk.$ (8)
[…]
@@
(See Sec. II of the Supplemental Material [2]) A net drift, characterized by $\langle v_{g}\rangle\neq 0$, emerges from the interplay between the initial state’s delocalization ($D$) and the Hamiltonian phase ($\alpha$). The drift vanishes for the extreme cases $D=0$ and $D=1$ but is maximized for intermediate delocalization, as shown in Fig. 2.
Figure 3: Time evolution of the mean square displacement (MSD) [Eq. (12)]. Solid lines represent the exact analytical results from Eq. (12), while symbols show numerical simulations. The top panel ($\alpha=0$) shows the quantum backfire effect: (i) for $t<t_{\mathrm{cross}}$, a larger initial MSD(0) promotes a larger MSD($t$); (ii) for $t>t_{\mathrm{cross}}$, this relationship inverts: a larger initial MSD(0) produces a smaller MSD($t$). The dashed vertical line marks $t_{\mathrm{cross}}$. The bottom panel ($\alpha=\pi/2$) exhibits no-crossing behavior: The ordering of MSD curves with respect to $D$ is preserved for all time.
Figure 4: Dependence of the MSD crossing time $t_{\mathrm{cross}}$ on the phase $\alpha$ [Eq. (13)]. Two regimes are shown. 1) No-crossing (yellow, $\sin^{2}\alpha\geq 1/2$): MSD curves remain ordered for all time. 2) Crossing (blue, $\sin^{2}\alpha<1/2$): MSD curves intersect at $t=t_{\mathrm{cross}}$. For $t>t_{\mathrm{cross}}$, the ordering inverts, demonstrating the quantum backfire effect where a greater initial delocalization ($D$) boosts short-time spreading, but is detrimental to the long-time propagation.
@@
The mean square displacement (MSD), defined as $\mathrm{MSD}(t)=\langle(X-X_{0})^{2}\rangle(t)$, measures the wavepacket’s spreading around its initial average position $X_{0}\equiv\langle X\rangle(0)$. For our initial state, symmetry dictates $\langle X\rangle(0)=0$, simplifying the expression to $\mathrm{MSD}(t)=\langle X^{2}\rangle(t)$. Evaluating the expectation value in the momentum basis, where $X=\mathrm{i}\partial_{k}$, allows us to decompose the MSD into a static initial contribution and a dynamic ballistic term (see Sec. III of the Supplemental Material for the detailed derivation [2]):
$\langle X^{2}\rangle(t)=\langle X^{2}\rangle(0)+t^{2}\langle v_{g}^{2}\rangle.$ (10)
The initial mean square position quantifies the initial delocalization: $\langle X^{2}\rangle(0)=D$. This result is intuitive, ranging from 0 for a localized state ($D=0$) to 1 for a state delocalized over sites $x=\pm 1$ ($D=1$).
The time-dependent spreading is governed by the expectation value of the squared group velocity, $\langle v_{g}^{2}\rangle$. Substituting $v_{g}(k)=-2\gamma\sin(\alpha-k)$ and the initial momentum distribution into the integral $\langle v_{g}^{2}\rangle=\int v_{g}(k)^{2}|\psi(k,0)|^{2}dk$ yields:
$\langle v_{g}^{2}\rangle=\gamma^{2}\left(2-D+2D\sin^{2}\alpha\right).$ (11)
The full step-by-step evaluation of this integral is provided in Sec. III of the Supplemental Material [2].
Combining these results yields the closed-form formula for the MSD:
$\mathrm{MSD}(t)=D+\gamma^{2}t^{2}\left(2-D+2D\sin^{2}\alpha\right).$ (12)
@@
$t_{\mathrm{cross}}(\alpha)=\frac{1}{\gamma\sqrt{\,1-2\sin^{2}\alpha\,}}.$ (13)
@@
The existence of a finite crossing time reveals a regime of counterintuitive dynamics. For a fixed time $t>t_{\mathrm{cross}}(\alpha)$ within the interval $0<\alpha<\pi/4$, Eq. 12 shows that the MSD becomes a decreasing function of the delocalization parameter $D$. This means that for these times, a more delocalized initial state (higher $D$) results in a smaller spatial spread of the wavepacket, the opposite of the intuitively expected outcome. This counterintuitive phenomenon, where the strategy of increasing initial delocalization to enhance spreading instead weakens it after a characteristic time, is formally analogous to the “Backfire Effect” known in cognitive science [40]. There, corrective evidence can paradoxically strengthen a person’s misconceptions. Here, a more delocalized initial state improves the short-time spreading, but leads to worse asymptotic spreading performance for a proper combination of parameters. We therefore designate this as a Quantum Backfire Effect.
This effect is clearly visible in the post-crossing regime (blue shading in Fig. 4 for $\sin^{2}\alpha<1/2$), where the MSD for $D=1$ becomes lower than for $D=0.5$, which in turn becomes lower than for $D=0$. For times $t<t_{\mathrm{cross}}(\alpha)$, the intuitive ordering, where greater initial delocalization leads to greater MSD, still holds.
Figure 5: Time evolution of $P_{\text{surv}}$ [Eq. (15)] on a log-log scale for $\alpha=\pi/2$. The $D=1$ case exhibits enhanced decay ($\sim t^{-3}$), while any partially delocalized state ($D<1$) shows the standard scaling ($\sim t^{-1}$). Dashed lines show the analytical asymptotic predictions.
@@
Starting with an initial condition where $P_{\text{surv}}(0)=1$, the wavepacket spreading leads to a decrease in this quantity. The decay analysis of this value informs how fast the particle leaves the region. Using the probability distribution we get (See Sec. IV of the Supplemental Material [2]):
@@
This exact result allows us to determine the asymptotic temporal decay by analyzing the envelope of the squared Bessel functions. For generic parameters, the survival probability follows the standard ballistic decay $P_{\text{surv}}(t)\propto t^{-1}$. However, for the fine-tuned case $D=1$ and $\alpha=\pi/2+m\pi$ ($m\in\mathbb{Z}$), the leading-order ($t^{-1}$) terms vanish due to destructive interference. In this regime, the decay is governed by the next order in the asymptotic expansion, resulting in $P_{\text{surv}}(t)\approx 1/(\pi\gamma^{3}t^{3})$.
These analytical results show that the enhanced temporal decay $t^{-3}$ occurs only in the case $D=1$ for specific values of $\alpha$. Other combinations of $D$ and $\alpha$ lead to the usual scaling with $t^{-1}$. These findings are numerically confirmed in Figure 5.
Final remarks
We presented an analytical treatment for CTQWs with a tunable delocalization parameter $D$. Our main results are given by Eqs. (9), (12), and (15), which relate the system’s transport properties to experimentally accessible observables. The key insights from our work are threefold.
First, we demonstrated the emergence of directional spreading from unbiased initial states with intermediate delocalization ($0<D<1$), resulting from the synergy between the initial state’s delocalization and the Hamiltonian’s phase.
Second, our analysis revealed a Quantum Backfire Effect: a phenomenon where increasing the initial state’s delocalization enhances short-time spreading but decreases the long-term spatial spread after a crossing time, $t_{\mathrm{cross}}$.
Third, by computing the exact survival probability, we elucidated that the transition from a standard scaling $\sim t^{-1}$ to an enhanced decay $\sim t^{-3}$ is a fine-tuned effect specific to the fully extended initial state ($D=1$) with specific phases.
The directional spreading and the backfire effect were not reported in related prior work [12]. Furthermore, we clarify that the breakdown of the standard scaling $P_{\mathrm{surv}}(t)\sim t^{-1}$ observed in [12] was a result of the specific initial condition chosen.
The control of transport dynamics in QWs is an important topic in quantum information science [36, 37, 24, 46, 25, 23, 27, 60, 34, 59, 51, 64, 65]. Our theoretical findings are accessible to current quantum simulation platforms [61]. For instance, photonic waveguide arrays allow for the engineering of complex hopping amplitudes via artificial gauge potentials [52, 31] and the precise preparation of delocalized initial states [56, 7], making the observation of these effects feasible.